The pressure of a sample of argon gas was increased from 1.36 atm to 6.20 atm at constant temperature. If the final volume of the argon sample was 11.6 L, what was the initial volume of the argon sample? Assume ideal behavior. V=___. L
Answer #1

Answer

by using Boyle's law
P1V1=P2V2
P1 = initial pressure = 1.36 atm
P2= final pressure = 6.20 atm
V1= initial volume= ?
V2= final volume = 11.6 L

V1= P2V2/P1 = (6.20*11.6)/1.36= 52.9 L